OpenStudy (ja1):

Question Below mate:

OpenStudy (ja1):

Where does the graph of \[y = 2x^{2} + 7x + 3\] cross the x-axis? I know this but I forgot it, need a refresher!

OpenStudy (ja1):

"WILL FAN AND MEDAL!!!" xD

terenzreignz (terenzreignz):

Just for that ^ I'm ignoring this :3

OpenStudy (ja1):

:( *sobs* ok....

terenzreignz (terenzreignz):

LOL Just kidding :)

OpenStudy (anonymous):

ahaha

OpenStudy (ja1):

I know ;)

terenzreignz (terenzreignz):

Cross the x-axis when y = 0 So... \[\large 0 = 2x^2 + 7x + 3\]

OpenStudy (anonymous):

set y=0

OpenStudy (ja1):

Gotcha...

OpenStudy (anonymous):

u do it

OpenStudy (ja1):

So which method do we use here?

terenzreignz (terenzreignz):

Using whatever method feels comfortable for you :)

OpenStudy (anonymous):

u need to get to 100 LOL

terenzreignz (terenzreignz):

Factor. 0 = (2x + 1)(x + 3)

OpenStudy (ja1):

Ah ok factoring is better for me to understand.

terenzreignz (terenzreignz):

Or you could just use the quadratic formula if you're feeling patient :D \[\Large x = \frac{-\color{blue}b\pm \sqrt{\color{blue}b^2 - 4\color{red}a\color{green}c}}{2\color{red}a}\]

OpenStudy (ja1):

Yes I think I will, I got confused just now D: One sec..

OpenStudy (ja1):

\[\frac{ -7 \pm \sqrt7^{2} - 4 (2) (3) }{ 2(2) }\]

OpenStudy (ja1):

Right?

OpenStudy (ja1):

Oops that square root got cut off xD

terenzreignz (terenzreignz):

\[\frac{ -7 \pm \sqrt{7^{2} - 4 (2) (3)} }{ 2(2) }\]

terenzreignz (terenzreignz):

Simplify? :D

OpenStudy (ja1):

Ok so it's right, good ok so here I go: \[\frac{ -7 \pm \sqrt{49 - 24} }{ 4 }\]

terenzreignz (terenzreignz):

Yup :) Continue... I'm watching you >:)

OpenStudy (ja1):

\[\frac{ -7 \pm \sqrt{25} }{ 4 }\] *Gulps* Oh Noez.. RUNNNN :D

terenzreignz (terenzreignz):

You're certainly one to take things step by step :/ \[\large \sqrt{25}=\color{red}?\]

OpenStudy (ja1):

\[\frac{ -7 + 5 }{ 4 }\] \[\frac{ -7 - 5 }{ 4 }\]

OpenStudy (ja1):

Yes I do xD but when it's in my head I do diferently.

terenzreignz (terenzreignz):

Okay... these aren't quite done yet :D

OpenStudy (ja1):

\[\frac{ -2 }{ 4 }\] \[\frac{ -12 }{ 4 }\]

terenzreignz (terenzreignz):

The suspense... it's killing...something :D Finish it up, mate :)

OpenStudy (ja1):

\[x = -0.5\] \[x = -3\]

OpenStudy (ja1):

LOL will do ;)

terenzreignz (terenzreignz):

Jawohl :D Just pointing out that these agree with the factored forms... \[0 = (2x+1)(x + 3)\] Implying 0 = 2x + 1 or 0 = x + 3 Would've saved you a few minutes if you'd just factored :D

OpenStudy (ja1):

Well I could of done the very minimal amount of quadratic equation but I like to do it thoroughly :3

terenzreignz (terenzreignz):

Well, good job, anyway ^_^

OpenStudy (ja1):

Thanks :D btw how did you get the factored version of this, that's why I was confused :/

OpenStudy (ja1):

You did quite wonderfully yourself sir ;)